Another outage somewhere in Asia ?

Anyone else noticed some major latency / packet loss going to Asia starting roughly 10 hours ago ?

For example, tracerouting to 36.68.218.127 provides very mixed results.

Traceroute from 192.168.121.126 to 36.68.218.127 (36.68.218.127):
1 192.168.121.1 1.337ms 0.375ms 0.34ms
2 lo0-0.bng4.wh-man.zen.net.uk (51.148.77.131) 5.744ms 4.817ms 4.139ms
3 lag-8.p2.wh-man.zen.net.uk (51.148.73.118) 11.71ms 11.515ms 11.817ms
4 lag-3.p1.thn-lon.zen.net.uk (51.148.73.136) 11.844ms 11.645ms 11.63ms
5 ldn-bb1-link.ip.twelve99.net (62.115.120.74) 12.61ms 12.755ms 12.758ms
6 prs-bb1-link.ip.twelve99.net (62.115.135.25) 19.242ms 18.703ms *
7 mei-b5-link.ip.twelve99.net (62.115.124.55) 28.8ms 28.702ms 28.593ms
8 * prs-bb1-link.ip.twelve99.net (62.115.124.54) 563.155ms *
9 * ldn-bb1-link.ip.twelve99.net (62.115.135.24) 637.818ms *
10 nyk-bb2-link.ip.twelve99.net (62.115.113.20) 364.331ms * *
11 * * *
12 port-b3-link.ip.twelve99.net (62.115.115.25) 1130.258ms 1155.178ms 1149.024ms
13 * tky-b3-link.ip.twelve99.net (62.115.127.11) 1109.197ms *
14 * tky-b2-link.ip.twelve99.net (62.115.126.134) 1498.216ms *
15 * * *
16 * * *
17 * * *
18 180.240.191.166 901.364ms * *
19 * 180.240.204.9 1164.257ms *
20 * * *
21 * * *
22 * * *
23 * * *
24 * * *
255 * * *

I assume this is due to cogent having massive outage in California and traffic shifting away from Cogent congesting 1299. It’s my quick observation

Oh, ok. That explains why it stops in Lax within Cogent. It appears to be a submarine cable cut from their page:

Customers Transiting our network from Europe to our Asian markets may be experiencing Latency and Packet Loss. Multiple vessels are currently being gathered and/or being dispatched to carry out the repair operations from the submarine faults currently being reported. There is NO ETR at this time.

This is within LA metro

We got bit by this overnight.

Customers transiting our network on the west coast of North America may have experienced Latency and Packet Loss. This was due to a fiber cut in our vendors network around the Los Angeles area. The vendor has found the damaged fiber and have repaired the cables. Cogent NOC is now monitoring this event. Customers should have seen their services restored at 11:46 AM ET. The master case is HD303782528.